OpenStudy (anonymous):

***Fan And Medal*** Four years ago, the water in the apartment complex pool was emptied using a single pump. That took 33 hours. To get the job done more quickly this year, the apartment manager is going to use an additional pump that would take 22 hours to empty the entire pool on its own. If the pool contains 99,000 gallons of water, how much faster will the pool empty using both pumps at the same time than it did four years ago?

OpenStudy (dan815):

hmm

OpenStudy (dan815):

okay so

OpenStudy (dan815):

add the work they both do together

OpenStudy (anonymous):

55 hours?

OpenStudy (dan815):

A--33 hhours to empty 99000 B--22 hours to empty 99000 gallons work of A = 99000/33 work of B=99000/22 99000/work of (A +B) = new time

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nice no wonder your a mathlete dan im an english girl total BOOKWORM!!!!

OpenStudy (dan815):

bahahah

OpenStudy (dan815):

that gave me 13+1/5 or 13.2

OpenStudy (anonymous):

Thank you!